OpenStudy (anonymous):

Solve for x? 2x + 1 = 5 @StudyGurl14

OpenStudy (anonymous):

2

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

2 times 2 is 4 + 1 = 5

OpenStudy (studygurl14):

You want to isolate the x value. First step: moving the 1 over to the right side You can do that by doing the opposite of 1, which is -1 2x + 1 = 5 -1 -1 ___________ 2x = 4 Next step: do the opposite of multiplying by two. What's the opposite of multiplying?

OpenStudy (anonymous):

oh i see, thank you @StudyGurl14 @Shroomsy You'd didive 2x by 2, and 4 by 2, to get x = 2

OpenStudy (studygurl14):

Yep. Good job. :)